OpenStudy (anonymous):

A flagpole is secured on opposite sides by two guy wires, each of which is 7 ft longer than the pole. The distance between the points where the wires are fixed to the ground is equal to the length of one guy wire. How tall is the flagpole (to the nearest inch)? ft in

OpenStudy (anonymous):

what is ____ft ? and ____in?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

45 ft, if I remember my formula correctly :) and 540 in

OpenStudy (anonymous):

Let p be the length of the pole. Solve the following for p:\[\text{Cos}\left[\text{ArcSin}\left[\frac{\frac{p+7}{2}}{p+7}\right]\right]= \frac{p}{p+7}\]\[p=\frac{7 \sqrt{3}}{2-\sqrt{3}}=45.2487 \]

OpenStudy (anonymous):

Trivial, but still...this is assuming the guy wires lie in the vertical plane, they are not slanted.